# Enhancing Dynamic CT Image Reconstruction with Neural Fields and Optical Flow

### v2

@@ 1.1 Dynamic Inverse Problems
For dynamic inverse problems, we formulate the problem as follows: we let the measurements $(f_{t})_{t\in\{t_{1},\ldots,t_{N_{T}}\}}\subset\mathbb{R}^{M}$, with $M$ the number of measurements collected at a given time. The goal is to recover a time-dependent quantity $u:\Omega_{T}:=\Omega\times[0,T]\to\mathbb{R}$ by solving the equation below
@@
Another option is explicitly considering the target’s motion by including its velocity field $\mathbb{v}$ in the formulation. We call this a motion-based regularizer, where a partial differential equation (PDE) $r(u,\mathbb{v})=0$, relating $u$, the velocity field $\mathbb{v}$, and their derivatives, is used to impose a physical prior. Usual examples include the optical flow and the continuity equations. In this paper, we focus on the former:
@@ 1.2 Neural Fields
In recent years, coordinate-based multilayer perceptrons (MLPs) have been employed as a new way of parameterizing quantities of interest. In computer vision, these are referred to as neural fields or implicit neural representations [13, 14], while the term Physics-Informed Neural Networks (PINNs) has been adopted when used to solve PDEs [15, 16]. The main idea is to use a neural network $u_{\theta}$ with trainable weights $\theta$ as an ansatz for the solution of the problem. It takes as input a spatio-temporal point $(x,t)\in\Omega_{T}$, and outputs the value $u_{\theta}(x,t)$ at that point, e.g., the intensity of the image at that particular time and location. The problem is then rephrased as a non-convex optimization that seeks optimal weights $\theta$. The method requires training a neural network for every new instance, thus, it is said to be self-supervised and differs from the usual learning framework where a solution map is found by training a network over large datasets. Applications of neural fields include image reconstruction in CT [17, 18, 19], MRI [20, 21, 22, 23, 24], image registration [25, 26, 27], continuous shape representation via signed distance functions [28], view synthesis [29], among others.
@@ 1.3 Contributions
In the context of dynamic inverse problems and neural fields, most of the literature relies entirely on the smoothness introduced by the network on the spatial and temporal variables to get a regularized solution. This motivates the minimization of the data-fidelity term without considering any explicit regularizers. Examples can be found on dynamic cardiac MRI in [21, 24], where the network outputs the real and imaginary parts of the signal, while in [22] the neural field is used to fit the measurements directly and then inference is performed by rendering the neural field at a uniform grid in the k-space followed by the inverse Fourier transform. In [38, 39] neural fields are used to solve a photoacoustic tomography dynamic reconstruction emphasizing their memory efficiency. Dynamic CT has been addressed in 3D+time domains in [19, 40], where the neural field parametrizes the initial frame and a deformation vector field warps it to get the subsequent frames in time. To the best of our knowledge, the previous are the only works using neural fields and a motion model via a deformable template.
Motivated by [8, 10] and leveraging the arbitrary resolution and automatic differentiation of neural fields to compute spatial and temporal derivatives, in this paper, we study neural fields in the context of dynamic inverse problems in a highly undersampled measurement regime with the optical flow equation as an explicit PDE-based motion regularizer imposed as a soft constraint as in PINNs. We do not consider the nuclear norm and the sparsity-based regularizers previously mentioned since they act on a discrete representation of the solution on a cartesian grid and hence do not exploit the mesh-free nature of the neural field: to use them, the neural field would need to be queried at points on a cartesian grid to get a discrete representation over which the regularizer can act. This has been done, for instance, in [23] for dynamic MRI.
Our findings are based on numerical experiments on dynamic CT performed on two synthetic and one real datasets, all in a 2D+time domain. The contributions are summarized as follows:
@@ 2 Dynamic Inverse Problems for Computed Tomography
Figure 2: Top row: ground truth image $u$ at frames 35 and 55 (out of 100) with two squares moving, and measurements $f\in\mathbb{R}^{M\times N_{T}}$. The red boxes indicate the projection obtained in frames 35 and 55. Bottom row: naive reconstructions using filtered back projection. The first two figures represent a frame-by-frame solution from measurements at frames 35 and 55. The third figure represents the static reconstruction from $f_{\text{bin}}$.
@@ 2.1 Motion Model
A motion model describes the relation between pixel intensities $u$ and the velocity flow $\mathbb{v}$ through an equation $r(u,\mathbb{v})=0$ in $\Omega_{T}$. Its choice is application-dependent, for instance, the continuity equation imposes a mass-preservation constraint, while the optical flow equation promotes no change in the intensities. These models are typically employed for the task of motion estimation, this is, determining the velocity flow $\mathbb{v}$ for the given image sequence $u$.
@@ 2.2 Joint Image Reconstruction and Motion Estimation
$\min_{u,\mathbb{v}}$ $\mathcal{D}(u,f_{t_{1}},\ldots,f_{t_{N_{T}}})+$ (5)
$\alpha\mathcal{R}(u)+\beta\mathcal{S}(\mathbb{v})+\gamma\mathcal{A}(r(u,\mathbb{v})),$
with $\alpha,\beta,\gamma>0$ being regularization parameters balancing the four terms. In [8], it is shown, among other things, how the pure motion estimation task of a noisy sequence can be enhanced by solving the joint task of image denoising and motion estimation.
This model was further employed for 2D+time problems in [10] and [11]. In the former, its application on dynamic CT is studied with sparse limited angles using both the $L^{1}$ and $L^{2}$-norms for the data fidelity term, with better results for the $L^{1}$-norm. In the latter, the same logic is used for dynamic cardiac MRI. In 3D+time domains, we mention [47] and [48] for dynamic CT and dynamic photoacoustic tomography respectively.
@@ 3 Methods
Depending on the nature of the noise, different data-fidelity terms can be considered. In this work, we consider Gaussian noise $\varepsilon$. To satisfy equation (2) at time $t$ we use an $L^{2}$ distance between predicted measurement and data $f$:
$\mathcal{D}_{t}(u,f):=\dfrac{1}{2}\|K_{t}[u_{t}]-f\|^{2}_{2}.$
@@
$\mathcal{D}(u,f_{t_{1}},\ldots,f_{t_{N_{T}}}):=\dfrac{1}{N_{T}}\displaystyle\sum_{i=1}^{N_{T}}\mathcal{D}_{t_{i}}(u,f_{t_{i}}).$ (6)
@@ 3.1 Numerical evaluation with Neural Fields
We parametrize both the image and the motion with two independent neural fields $u_{\theta}$ and $\mathbb{v}_{\phi}$, respectively, with $\theta$ and $\phi$ denoting the weights and biases of the networks. Even though $u_{\theta}$ is a continuous and explicit representation, there is not a closed-form expression for its X-ray transform $K_{t}[(u_{\theta})_{t}]$. For this reason, the numerical evaluation of the forward operator is performed by first evaluating the network at points on a cartesian grid to get a grid-based representation at time $t$ as $(u_{\theta})_{t}:=\{u_{\theta}(x_{j},t)\}_{j=1,\ldots,N}$, and then applying our preferred implementation of the X-ray transform. In particular, we work with Tomosipo [49], a library that provides an integration of the ASTRA-toolbox [50, 51] with PyTorch, making it suitable for the optimization of neural networks. The data fidelity term in (6) requires first the evaluation of the network at $N\times N_{T}$ fixed grid-points to get the scene $\{u_{\theta}(x_{j},t_{i})\}_{i=1,\ldots,N_{T};j=1,\ldots,N}$, and, second, the application of the forward models $\{K_{t_{i}}\}_{i=1\ldots,N_{T}}$ to each frame. We call this the full-batch approach. This might be expensive and time-consuming. A common practice is to proceed with a mini-batch-like approach in time to speed up the optimization and avoid poor local minima. In this setting, at each iteration, we randomly sample $1\leq N_{b}\leq N_{T}$ frames, say, $\{i_{1},\ldots,i_{N_{b}}\}$, and the neural field is evaluated at the corresponding points to get the representation of the image at times $\{t_{i_{1}},\ldots,t_{i_{N_{b}}}\}$. Then, the forward model is applied on these frames only and the parameters are updated to minimize the difference between predicted data and the measured data $\{f_{t_{i_{1}}},\ldots,f_{t_{i_{N_{b}}}}\}$. This represents considerable benefits in terms of memory since the whole scene is never explicitly represented in the whole space-time grid and the cost of applying the forward operator on many frames is avoided, however, it adds variability during optimization.
@@
Since neural fields are mesh-free, the regularization terms can be evaluated at any point of the domain. Additionally, derivatives can be computed through automatic differentiation. This motivates approximating the integrals for the regularizers in equation (7) via Monte Carlo by randomly sampling $N_{c}$ collocation points of the form $\{(x_{c},t_{c})\}_{c=1,\ldots,N_{c}}\subset\Omega_{T}$ using a Latin Hypercube Sampling strategy [52]. For ease of notation, we introduce the function $\eta$ defined as the integrand in the regularization terms:
@@
Finally, we recall that it is an open question how to choose $N_{c}$, the number of collocation points sampled at each iteration. One would like to sample as many points as possible to have a better approximation of the regularizer, however, this might be time-consuming and prohibitive in terms of memory because of the use of auto differentiation. Thus, we define the sampling rate ($SR$) as the ratio between $N_{c}$ and the amount of points $N_{T}\times N$ on the spatiotemporal grid:
@@ 3.2 Numerical evaluation with grid-based representation
¯$\overline{\mathcal{R}}(u)$ $:=\dfrac{|\Omega_{T}|}{N_{T}N}\sum_{i=1}^{N_{T}}\|(Du)_{i}\|_{2,1}$
$\bar{\mathcal{A}}(r(u,v))$ $:=\dfrac{|\Omega_{T}|}{N_{T}N}\sum_{i=1}^{N_{T}}\|(D_{t}u)_{i}+\mathbb{v}_{i}\cdot(Du)_{i}\|_{1},$
@@
$\min_{u,v}$ $\mathcal{D}(u,f_{t_{1}},\ldots,f_{t_{N_{T}}})+$ (9)
$\alpha\overline{\mathcal{R}}(u)+\beta\overline{\mathcal{S}}(\mathbb{v})+\gamma\bar{\mathcal{A}}(r(u,\mathbb{v})),$
Due to the non-linearity present in the optical flow equation, this problem is not convex, however, it can be easily seen that it is biconvex, hence, in [8], the proposed optimization routine updates the current iteration $(u^{k},\mathbb{v}^{k})$ by alternating between the following two subproblems:
@@
    $u^{k+1}=\operatorname*{arg\,min}_{u}$ $\mathcal{D}(u,f_{t_{1}},\ldots,f_{t_{N_{T}}})+$ (10)
    $\alpha\overline{\mathcal{R}}(u)+\gamma\bar{\mathcal{A}}(r(u,\mathbb{v}^{k}))$
@@
    $\mathbb{v}^{k+1}=\operatorname*{arg\,min}_{v}$ $\beta\overline{\mathcal{S}}(\mathbb{v})+\gamma\bar{\mathcal{A}}(r(u^{k+1},\mathbb{v})).$ (11)
@@ 4 Datasets
We study our method in a 2D+time setting. We employ two synthetic datasets, the two-square and the cardiac phantoms, and a real one, the STEMPO phantom [54]. For all the experiments the considered physical domain is the square $\Omega=[-1,1]^{2}$. The code will be made available on Github upon acceptance of the paper.
For the synthetic experiments, we define the ground-truth phantom $u:\Omega_{T}$ as follows:
@@
  • Define $\varphi:\Omega_{T}\to\Omega$ describing the motion of the process. It takes a point $(x_{0},y_{0},t)\in\Omega_{T}$ and outputs $\varphi(x_{0},y_{0},t)\in\Omega$, the new position of $(x_{0},y_{0})$ at time $t$. For each time we can define the function $\varphi_{t}:\Omega\to\Omega$ by $\varphi_{t}(x_{0},y_{0})=\varphi(x_{0},y_{0},t)$. We require $\varphi_{t}$ to be a diffeomorphism for every $t\in[0,T]$. Hence we can define the trajectory of the point $(x_{0},y_{0})$ as $t\to\varphi_{t}(x_{0},y_{0})$.
  • Define $u(x,y,t):=u_{0}(\varphi_{t}^{-1}(x,y))$.
The phantom $u$ generated by the above procedure solves the optical flow equation with velocity field $\mathbb{v}=\frac{d}{dt}\varphi$. To mimic the continuous world, we define the ground truth at a high spatial resolution of $1024\times 1024$. Measurements are generated using Tomosipo assuming a camera with $M=64$ sensors. To avoid the inverse crime, during the reconstruction, we evaluate the neural field at a lower resolution spatial grid to get the discretized image as described in section 3.1.
For the synthetic datasets, we consider two sampling strategies, random and sequential. For the first, at each time a projection is taken along a random angle $\theta\in[0,2\pi)$; for the second, projections are acquired at fixed 9-degree intervals between consecutive frames. For STEMPO, both strategies are sequential, the first one with 32-degree intervals and the second with 4-degree intervals between frames.

### v3

@@ 1.1 Dynamic Inverse Problems
For dynamic inverse problems, we formulate the problem as follows: we let $\mathbb{f}=\{f_{t_{1}},\ldots,f_{t_{N_{T}}}\}\subset\mathbb{R}^{M}$ be the measurements at several time steps, with $M$ the number of measurements collected at a given time. The goal is to recover a time-dependent quantity $u:\Omega_{T}:=\Omega\times[0,T]\to\mathbb{R}$ by solving the equation below
@@
Another option is explicitly considering the target’s motion by including its velocity field $\mathbb{v}$ in the formulation. We call this a motion-based regularizer, where a partial differential equation (PDE) $r(u,\mathbb{v})=0$, relating $u$, the velocity field $\mathbb{v}$, and their derivatives, is used to impose a physical prior. In this paper, we set the motion model as the so-called Optical Flow equation:
@@ 1.2 Neural Fields
In recent years, coordinate-based multilayer perceptrons (MLPs) have been employed as a new way of parameterizing quantities of interest. In computer vision, these are referred to as neural fields or implicit neural representations [13, 14], while the term Physics-Informed Neural Networks (PINNs) has been adopted when used to solve PDEs [15, 16]. The main idea is to use a neural network $u_{\theta}$ with trainable weights $\theta$ as an ansatz for the solution of the problem. It takes as input a spatio-temporal point $(x,t)\in\Omega_{T}$, and outputs the value $u_{\theta}(x,t)$ at that point, e.g., the intensity of the image at that particular time and location. The problem is then rephrased as a non-convex optimization that seeks optimal weights $\theta$. The method requires training a neural network for every new instance, thus, it is said to be self-supervised and differs from the usual learning framework where a solution map is found by training a network over large datasets. Applications of neural fields include image reconstruction in CT [17, 18, 19], MRI [20, 21, 22, 23, 24], image registration [25, 26, 27], continuous shape representation via signed distance functions [28], view synthesis with Neural Radiance Fields (NeRF) [29], among others.
@@
In the context of dynamic inverse problems and neural fields, part of the literature relies entirely on the smoothness introduced by the network on the spatial and temporal variables to get a regularized solution. This motivates the minimization of the data-fidelity term without considering any explicit regularizers. Examples can be found on dynamic cardiac MRI in [21, 24], where the network outputs the real and imaginary parts of the signal. In [17, 18, 38] a neural field is used to inpaint the undersampled sinogram. Once optimized, inference is performed by rendering the network at a regular grid and applying a suitable transformation, e.g., filtered back projection. Regularized variational problems with neural fields have been considered in [22, 39] with a regularization-by-denoising approach for MRI and intensity diffraction tomography respectively, while [20, 23] approximate a total variation regularizer with finite differences. Such approaches do not exploit the continuous and differentiable representation offered by neural fields. In [40, 41] neural fields are used to solve a photoacoustic tomography dynamic reconstruction emphasizing their memory efficiency and using total variation and Tikhonov regularization respectively with gradients computed via automatic differentiation. Dynamic CT has been addressed in 3D+time domains in [19, 42], where the neural field parametrizes the initial frame and a deformation vector field warps it to get the subsequent frames in time. A similar idea is used for novel view synthesis for dynamic scenes in D-NeRF [43].
@@ 1.3 Contributions
Motivated by [8, 10], in this paper, we study neural fields in the context of dynamic inverse problems in a highly undersampled measurement regime with the optical flow equation as an explicit PDE-based motion regularizer imposed as a soft constraint as in PINNs. We leverage the arbitrary resolution and automatic differentiation of neural fields to compute spatial and temporal derivatives. We do not consider the nuclear norm and the sparsity-based regularizers previously mentioned since they act on a discrete representation of the solution on a cartesian grid and hence do not exploit the mesh-free nature of the neural field: to use them, the neural field needs to be queried at points on a cartesian grid to get a discrete representation over which the regularizer can act.
Our findings are based on numerical experiments on dynamic CT performed on three synthetic and one real datasets, all in a 2D+time domain. The contributions are summarized as follows:
@@ 2 Dynamic Inverse Problems for Computed Tomography
Figure 2: Top row: ground truth image $u$ at frames 35 and 55 (out of 100) with two squares moving, and measurements $\mathbf{f}\in\mathbb{R}^{M\times N_{T}}$. The red boxes indicate the projection obtained in frames 35 and 55. Bottom row: naive reconstructions using filtered back projection. The first two figures represent a frame-by-frame solution from measurements at frames 35 and 55. The third figure represents the static reconstruction from $f_{\text{bin}}$.
@@ 2.1 Motion Model
A motion model describes the relation between pixel intensities $u$ and the velocity flow $\mathbb{v}$ through an equation $r(u,\mathbb{v})=0$ in $\Omega_{T}$. See, e.g., equation (3). Its choice is application-dependent, for instance, the continuity equation imposes a mass-preservation constraint, while the optical flow equation promotes no change in the intensities. These models are typically employed for the task of motion estimation, this is, determining the velocity flow $\mathbb{v}$ for the given image sequence $u$.
@@ 2.2 Joint Image Reconstruction and Motion Estimation
$\min_{u,\mathbb{v}}$ $\mathcal{D}(u,\mathbb{f})+\alpha\mathcal{R}(u)+\beta\mathcal{S}(\mathbb{v})+\gamma\mathcal{A}(r(u,\mathbb{v})),$ (5)
where $\alpha,\beta,\gamma>0$ are regularization parameters balancing the four terms. We also recall $\mathbb{f}=\{f_{t_{1}},\ldots,f_{t_{N_{T}}}\}$. In [8], it is shown, among other things, how the pure motion estimation task of a noisy sequence can be enhanced by solving the joint task of image denoising and motion estimation.
This model was further employed for 2D+time problems in [10] and [11]. In the former, its application on dynamic CT is studied with sparse limited angles using both the $L^{1}$ and $L^{2}$-norms for the data fidelity term, with better results for the $L^{1}$-norm. In the latter, the same logic is used for dynamic cardiac MRI. In 3D+time domains, we mention [50] and [51] for dynamic CT and dynamic photoacoustic tomography respectively. More methods for dynamic CT have been proposed based on the simultaneous algebraic reconstruction technique and motion compensation. We refer to the interested reader to [52, 53]
@@ 3 Methods
Different data-fidelity terms can be considered depending on the nature of the noise. In this work, we consider Gaussian noise $\varepsilon$. To satisfy equation (2) at time $t$ we use an $L^{2}$ distance between predicted measurement and data $f_{t}$:
$\mathcal{D}_{t}(u,f_{t}):=\dfrac{1}{2}\|K_{t}[u_{t}]-f_{t}\|^{2}_{2}.$
@@
$\mathcal{D}(u,\mathbb{f}):=\dfrac{1}{N_{T}}\displaystyle\sum_{i=1}^{N_{T}}\mathcal{D}_{t_{i}}(u,f_{t_{i}}).$ (6)
@@ 3.1 Numerical evaluation with Neural Fields
We parametrize both the image and the motion with two independent neural fields. Both take a point $(x,t)\in\mathbb{R}^{3}$ as input, then a Fourier feature embedding [34] is applied independently on the space and time variables, and then apply $L+1$ fully-connected transformations. This is described below:
$\boldsymbol{x}_{0}$ $=(\Gamma_{1}(x),\Gamma_{2}(t))\in\mathbb{R}^{m}$ (8)
$\boldsymbol{x}_{l}$ $=\sigma(W^{l}\boldsymbol{x}_{l-1}+b^{l})\in\mathbb{R}^{d_{l}},\quad l=1,\ldots,L,$
$\boldsymbol{x}_{L+1}$ $=W^{L+1}\boldsymbol{x}_{L}+b^{L+1}\in\mathbb{R}^{d_{L+1}},$
where $\{(W^{l},b^{l})\}_{l=1}^{L+1}$ are the weights and biases and $\sigma:\mathbb{R}\to\mathbb{R}$ is the non-linear activation function acting element-wise. The Fourier embeddings are defined as $\Gamma_{1}(x):=(\sin(2\pi\mathbf{B_{x}}x),\cos(2\pi\mathbf{B_{x}}x))\in\mathbb{R}^{2m_{x}}$ and $\Gamma_{2}(t):=(\sin(2\pi\mathbf{B_{t}}t),\cos(2\pi\mathbf{B_{t}}t))\in\mathbb{R}^{2m_{t}}$, with the sinusoidal functions acting element-wise. The matrices $\mathbf{B_{x}}\in\mathbb{R}^{m_{x}\times 2}$ and $\mathbf{B_{t}}\in\mathbb{R}^{m_{t}\times 1}$ have non-trainable entries sampled from gaussian distributions $(\mathbf{B_{x}})_{ij}\sim\mathcal{N}(0,\sigma_{x}^{2})$ and $(\mathbf{B_{t}})_{ij}\sim\mathcal{N}(0,\sigma_{t}^{2})$. Here $\sigma_{x}$ and $\sigma_{t}$ are hyperparameters accounting for the frequencies the neural field can capture; the larger they are, the more frequencies can be captured earlier during optimization.
@@
We let $u_{\theta}$ and $\mathbb{v}_{\phi}$ to be the image and the velocity field, respectively, with $\theta$ and $\phi$ denoting their weights and biases. Clearly, we set $d_{L+1}=1$ for $u_{\theta}$ and $d_{L+1}=2$ for $\mathbb{v}_{\phi}$. Even though $u_{\theta}$ is a composition of known functions, there is no closed-form expression for its X-ray transform $K_{t}[(u_{\theta})_{t}]$. For this reason, the numerical evaluation of the forward operator is performed by first evaluating the network at points on a cartesian grid to get a grid-based representation at time $t$ as $(u_{\theta})_{t}:=\{u_{\theta}(x_{j},t)\}_{j=1,\ldots,N}$, and then applying our preferred implementation of the X-ray transform. In particular, we work with Tomosipo [54], a library that provides an integration of the ASTRA-toolbox [55, 56] with PyTorch, making it suitable for the optimization of neural networks. The data fidelity term in (6) requires first the evaluation of the network at $N\times N_{T}$ fixed grid-points to get the scene $\{u_{\theta}(x_{j},t_{i})\}_{i=1,\ldots,N_{T};j=1,\ldots,N}$, and, second, the application of the forward models $\{K_{t_{i}}\}_{i=1\ldots,N_{T}}$ to each frame. We call this the full-batch approach. This might be expensive and time-consuming. A common practice is to proceed with a mini-batch-like approach in time to speed up the optimization and avoid poor local minima. In this setting, at each iteration, we randomly sample $1\leq N_{B}\leq N_{T}$ frames, say, $\{i_{1},\ldots,i_{N_{B}}\}$, and the neural field is evaluated at the corresponding points to get the representation of the image at times $\{t_{i_{1}},\ldots,t_{i_{N_{B}}}\}$. Then, the forward model is applied on these frames only and the parameters are updated to minimize the difference between predicted data and the measured data $\{f_{t_{i_{1}}},\ldots,f_{t_{i_{N_{B}}}}\}$. This represents considerable benefits in terms of memory since the whole scene is never explicitly represented in the whole space-time grid and the cost of applying the forward operator on many frames is avoided, however, it adds variability during optimization.
@@
Since neural fields are mesh-free, the regularization terms can be evaluated at any point of the domain. Additionally, derivatives can be computed through automatic differentiation. This motivates approximating the integrals for the regularizers in equation (7) via Monte Carlo by randomly sampling $N_{C}$ collocation points of the form $\{(x_{c},t_{c})\}_{c=1,\ldots,N_{C}}\subset\Omega_{T}$. For ease of notation, we introduce the function $\eta$ defined as the integrand in the regularization terms:
@@
Finally, we recall that it is an open question how to choose $N_{C}$, the number of collocation points sampled at each iteration. One would like to sample as many points as possible to have a better approximation of the regularizer, however, this might be time-consuming and prohibitive in terms of memory because of the use of auto differentiation. Thus, we define the sampling rate ($SR$) as the ratio between these collocation points and the amount of points on the spatiotemporal grid:
@@ 3.2 Numerical evaluation with grid-based representation
¯$\overline{\mathcal{R}}(u)$ $:=\dfrac{|\Omega_{T}|}{N_{T}N}\sum_{i=1}^{N_{T}}\|(Du)_{i}\|_{2,1},$
$\overline{\mathcal{A}}(r(u,\mathbb{v}))$ $:=\dfrac{|\Omega_{T}|}{N_{T}N}\sum_{i=1}^{N_{T}}\|(D_{t}u)_{i}+\mathbb{v}_{i}\cdot(Du)_{i}\|_{1},$
@@
$\min_{u,\mathbb{v}}$ $\mathcal{D}(u,\mathbb{f})+\alpha\overline{\mathcal{R}}(u)+\beta\overline{\mathcal{S}}(\mathbb{v})+\gamma\overline{\mathcal{A}}(r(u,\mathbb{v})),$ (10)
This problem is non-convex due to the non-linearity present in the optical flow equation, however, it can be easily seen that it is biconvex, hence, in [8], the proposed optimization routine updates the current iteration $(u^{k},\mathbb{v}^{k})$ by alternating between the following two subproblems:
@@
    $u^{k+1}=\operatorname*{arg\,min}_{u}$ $\mathcal{D}(u,\mathbb{f})+\alpha\overline{\mathcal{R}}(u)+\gamma\overline{\mathcal{A}}(r(u,\mathbb{v}^{k}))$ (11)
@@
    $\mathbb{v}^{k+1}=\operatorname*{arg\,min}_{\mathbb{v}}$ $\beta\overline{\mathcal{S}}(\mathbb{v})+\gamma\overline{\mathcal{A}}(r(u^{k+1},\mathbb{v}))$ (12)
@@ 4 Datasets
We study our method in a 2D+time setting. We employ three synthetic datasets, the two-square, the cardiac, and the XCAT phantoms, and a real one, the STEMPO phantom [58]. For all the experiments the considered physical domain is the square $\Omega=[-1,1]^{2}$. Code will be made available on Github upon acceptance of the paper.
For the two-square and cardiac phantoms, we define the ground-truth phantom $u:\Omega_{T}\to[0,1]$ as follows:
@@
  • Define $\mathbb{\varphi}:\Omega_{T}\to\Omega$ describing the motion of the process. It takes a point $(x_{0},y_{0},t)\in\Omega_{T}$ and outputs $\varphi(x_{0},y_{0},t)\in\Omega$, the new position of $(x_{0},y_{0})$ at time $t$. For each time we can define the function $\mathbb{\varphi}_{t}:\Omega\to\Omega$ by $\mathbb{\varphi}_{t}(x_{0},y_{0})=\mathbb{\varphi}(x_{0},y_{0},t)$. We require $\mathbb{\varphi}_{t}$ to be a diffeomorphism for every $t\in[0,T]$. Hence we can define the trajectory of the point $(x_{0},y_{0})$ as $t\to\mathbb{\varphi}_{t}(x_{0},y_{0})$.
  • Define $u(x,y,t):=u_{0}(\mathbb{\varphi}_{t}^{-1}(x,y))$.
The phantom $u$ generated by the above procedure solves the optical flow equation with velocity field $\mathbb{v}=\frac{d}{dt}\mathbb{\varphi}$. To mimic the continuous world, we define the ground truth at a high spatial resolution of $1024\times 1024$. Measurements are generated using Tomosipo, assuming a camera with $M=64$ sensors. To avoid the inverse crime, during the reconstruction, we evaluate the neural field at a lower resolution spatial grid to get the discretized image as described in section 3.1.
We shall consider two sampling strategies: random and sequential. For the first, at each frame a projection is taken along a random angle $\theta\in[0,2\pi)$; for the second, projections are acquired at fixed 9-degree intervals between consecutive frames. For STEMPO, both strategies are sequential, the first one with 32-degree intervals and the second with 4-degree intervals between frames.